# Integrating Large Language Models into a Tri-Modal Architecture for Automated Depression Classification

### v4

@@ III-C Text Preprocessing
0 | Ellie | how_doingV (so how are you doing today) |
1 | Participant | good |
2 | Participant | it’s been a nice day |
3 | Ellie | thats_good (that’s good) |
4 | Ellie | where_originally (where are you from originally) |
@@
0 | Therapist | So how are you doing today? |
1 | Patient | Good. It’s been a nice day. |
2 | Therapist | That’s good. Where are you from originally? |
@@ III-E Visual Preprocessing
1 | AU01_r | Inner Brow Raiser |
2 | AU02_r | Outer Brow Raiser |
3 | AU04_r | Brow Lowerer |
4 | AU05_r | Upper Lid Raiser |
5 | AU06_r | Cheek Raiser |
6 | AU09_r | Nose Wrinkler |
7 | AU10_r | Upper Lip Raiser |
8 | AU12_r | Lip Corner Puller |
9 | AU14_r | Dimpler |
10 | AU15_r | Lip Corner Depressor |
11 | AU17_r | Chin Raiser |
12 | AU20_r | Lip Stretcher |
13 | AU25_r | Lips Part |
14 | AU26_r | Jaw Drop |
15 | AU04_c | Brow Lowerer |
16 | AU12_c | Lip Corner Puller |
17 | AU15_c | Lip Corner Depressor |
18 | AU23_c | Lip Tightener |
19 | AU28_c | Lip Suck |
20 | AU45_c | Blink |

### v5

@@ III-C Text Preprocessing
| 0 | Ellie | how_doingV (so how are you doing today) |
| 1 | Participant | good |
| 2 | Participant | it’s been a nice day |
| 3 | Ellie | thats_good (that’s good) |
| 4 | Ellie | where_originally (where are you from originally) |
@@
| 0 | Therapist | So how are you doing today? |
| 1 | Patient | Good. It’s been a nice day. |
| 2 | Therapist | That’s good. Where are you from originally? |
@@ III-E Visual Preprocessing
| 1 | AU01_r | Inner Brow Raiser |
| 2 | AU02_r | Outer Brow Raiser |
| 3 | AU04_r | Brow Lowerer |
| 4 | AU05_r | Upper Lid Raiser |
| 5 | AU06_r | Cheek Raiser |
| 6 | AU09_r | Nose Wrinkler |
| 7 | AU10_r | Upper Lip Raiser |
| 8 | AU12_r | Lip Corner Puller |
| 9 | AU14_r | Dimpler |
| 10 | AU15_r | Lip Corner Depressor |
| 11 | AU17_r | Chin Raiser |
| 12 | AU20_r | Lip Stretcher |
| 13 | AU25_r | Lips Part |
| 14 | AU26_r | Jaw Drop |
| 15 | AU04_c | Brow Lowerer |
| 16 | AU12_c | Lip Corner Puller |
| 17 | AU15_c | Lip Corner Depressor |
| 18 | AU23_c | Lip Tightener |
| 19 | AU28_c | Lip Suck |
| 20 | AU45_c | Blink |